OpenStudy (anonymous):

Can someone provide some assistance please? Use the functions f(x) = 5x − 12 and g(x) = 11x + 2 to complete the function operations listed below. Part A: Find (f + g)(x). Show your work. (3 points) Part B: Find (f ⋅ g)(x). Show your work. (3 points) Part C: Find f[g(x)]. Show your work. (4 points)

OpenStudy (anonymous):

I believe that for the first part that you're supposed to add the two equations, but the part about part A that confuses me is /how/ do you add them?

OpenStudy (anonymous):

yes for part A you add. \[(f+g)(x)=f(x)+g(x)=5x-12+11x+2\]

OpenStudy (anonymous):

Combine like terms

OpenStudy (anonymous):

Would it be -6x?

OpenStudy (anonymous):

you'd add the terms with x's together and the constant terms, so (5x + 11x) + (-12 + 2) Do each grouping separately

OpenStudy (anonymous):

16x + (-10)?

OpenStudy (anonymous):

right, so 16x - 10 is the answer

OpenStudy (anonymous):

Oh okay! Can you help with the other two parts as well, please?

OpenStudy (anonymous):

yeah. For B you have to use FOIL \[(5x-12)(11x+2)\]

OpenStudy (anonymous):

\[55x^2-122x-24\] Right?

OpenStudy (anonymous):

yes that's right

OpenStudy (anonymous):

For C, they want f(g(x)), so you substitute g(x) for the x in f(x)